OpenStudy (anonymous):

The midpoint of a line segment with end points as (2, 3) and (4, y2) is (3, 2). What is the value of y2? Answer -2 -1 1 7

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1

OpenStudy (anonymous):

If you want to know why, ask

OpenStudy (anonymous):

thamk u and what confused on the last reply

OpenStudy (anonymous):

I was just asking if you wanted to know how we got this answer

OpenStudy (anonymous):

oh ya could u show me i dont really have a teacher im homeschooled

OpenStudy (anonymous):

ok. the midpoint of a line segment is just the average of each coordinate value. In this case: \[\frac{4+2}{2}=3\]for the first coordinate, and,\[\frac{y_{2}+3}{2}=2\]Solving for y2, you get y2=1

OpenStudy (anonymous):

ohhh okay thank u so much like i said i sux at geometry probably because i dont have a teacher :/

OpenStudy (anonymous):

good thing for open study, there's always competent folks to help out :)